OpenStudy (anonymous):

if tan (b) = squareroote (5)/ 2 what is cot(b) what is sin (b) and what is cos (b)

OpenStudy (anonymous):

tan = opp/adj cot = 1/tan sin = opp /hyp coss = adj / hyp tan = sin/cos

OpenStudy (anonymous):